OpenStudy (anonymous):

Any help you could give me with this problem would be grand! tan theta=-1 I already found the equation that expresses the value of theta in terms of an appropriate inverse trigonometric expression: theta=tan^-1(-1) But I need help answering: on the inverval [0, 2pi], what values of theta satisfy that equation? Would it be 3pi/4 and -pi/4. Also, how would you write an expression for all solutions to the equation?

OpenStudy (anonymous):

since tangent it periodic with period \(\pi\) youcoud say \[\frac{3\pi}{4}+k\pi\] for \(k\in \mathbb{Z}\)

OpenStudy (anonymous):

But I need help answering: on the inverval [0, 2pi], Would it be 3pi/4 and -pi/4. NO

OpenStudy (anonymous):

because \(-\frac{\pi}{4}\) is not in the interval \([0,2\pi)\)

OpenStudy (anonymous):

in that interval you would say \[\frac{3\pi}{4},\frac{7\pi}{4}\]

OpenStudy (anonymous):

thank you so much!!!! @satellite73